A solution to be tested by AAS is prepared by perf
Answer #1

Dilution law: M1V1 = M2V2

M1 and V1 are concentration and volume before dilution.

M2 and V2 are concentration and volume after dilution.

M1 x 0.10 mL = 3.4 ppm x 100.0 mL

M1 = ( 3.4 ppm x 100.0 mL)/ 0.10 mL = 3400 ppm

Concentration of Cd2+ in starting solution is 3400 ppm.

ppm = 1 micro gram/1 L

3400 ppm = 3400x10-6 g/1L

Atomic mass of Cd is 112 g/mol

So, (3400x10-6 g)/(112 g/mol)/1L = 3.036x10-5 mol/L = 3.036x10-5 M